OpenStudy (anonymous):

Find the number c that satisfies the conclusion of Rolle's Theorem. f(x) = 5 - 8x + 2x^2 [1, 3]

myininaya (myininaya):

f'(x)=-8+4x f'(x)=0 imples -8+4x=0 implies 4x=8 so x=2 since x=2 in between 1 and 3 then x=2 is the c that satisfies the conclusion of rolles thm

OpenStudy (anonymous):

thats it??

OpenStudy (anonymous):

Don't forget, for Rolle's theorem to be applicable, you need to show that f(1)=f(3) (which it does here).

OpenStudy (anonymous):

and what if it doesnt

myininaya (myininaya):

i assumed that was the case since it said just to look at the conclusion

OpenStudy (anonymous):

it does...i was just wondering

myininaya (myininaya):

but it easy to check to see if it does satisfy the hypothese so why not

OpenStudy (anonymous):

Rolle's theorem is dependent upon f real being continuous on a closed interval [a,b], differentiable on the open interval, (a,b), and f(a)=f(b).

myininaya (myininaya):

right vg

OpenStudy (anonymous):

You would roll over to the mean value theorem if f(a) did not equal f(b).

OpenStudy (anonymous):

ok i get it! thank u guys! :D

OpenStudy (anonymous):

kk

myininaya (myininaya):

but who needs rolle when u have mean value thm it easier to memorize one thm than two

OpenStudy (anonymous):

agreed :)